OpenStudy (anonymous):

PLEASE HELP!! Which of the following is the solution of log base x + 5 0.001 = –3 ? x = 5 x = 7 x = 13 x = 15

OpenStudy (anonymous):

\[\log _{x+5}0.001 = -3\]

OpenStudy (anonymous):

means you have to solve \((x+5)^{-3}=0.001\)

OpenStudy (anonymous):

and since \(10^{-3}=.001\) you know \(x+5=10\)

OpenStudy (cwrw238):

by laws of logarithms if log base a of x = y then a^y = x - as satellite says

OpenStudy (anonymous):

so x = 5

OpenStudy (cwrw238):

yes